# Non-Markovian Dynamics in Fiber Delay-Line Buffers

### v2

@@ IV Methods
Atom-Cavity-Reservoir Model. Our analytical approach is similar to a qubit of a two-level atom interacting in a high Q-cavity, which is coupled to the reservoir with the Lorentzian spectral distribution Chen_PRA_14 ; Wang_scr_17 . The cavity-reservoir coupling gives rise to the effective spectral density in the form of, $(\gamma_{\circ}\Lambda^{2})/[(\omega_{\circ}-\omega)^{2}+\Lambda^{2}]$, where the parameter $\gamma_{\circ}$ is the coupling constant between the cavity and the reservoir. The parameter $\Lambda$ is the spectral width of the reservoir. The atom transition frequency and the cavity are in resonance with $\omega_{\circ}=\omega_{\rm{cavity}}$. When the $\Lambda$ becomes very large or approaches infinity, the reservoir is considered memoryless or Markovian. Then, the cavity is only the main cause of the non-Markovian dynamic in the system. The diagonal (off-diagonal)density matrix elements of the atomic qubit are evolving in time $t$ as a function of $p(t)\propto e^{-\frac{\gamma_{\circ}t}{2}}[cos(\frac{\delta t}{4})+\frac{\gamma_{\circ}}{\delta}sin(\frac{\delta t}{4})]^{2}$ ($\sqrt{p(t)}$), respectively, where $\delta=\sqrt{16\kappa^{2}-\gamma_{\circ}^{2}}$, and $\kappa$ is the coupling constant between the atomic qubit and the cavity and related to the decay rate of the excited state. The parameter $\delta$ can be used to determine whether the system is Markovian ( $4\kappa<\gamma_{\circ}$) or non-Markovian ($4\kappa\geq\gamma_{\circ}$).
@@
For $4\kappa<\gamma_{\circ}$ or $1<\frac{\gamma_{\circ}}{4\kappa}$, $\delta\rightarrow i\delta$, the harmonic function $sin(\cdot)$ and $cos(\cdot)$ becomes hyperbolic function $cosh(\cdot)$ and $sinh(\cdot)$, respectively. Our quantum system is undergoing Markovian dynamics. This occurs when the PMD and the loss parameter are the dominant effects on the photonic qubit. Hence, the probability function $p(t)$ becomes an exponential function. The polarization mode dispersion (PMD) effect that leads to the Markovian process i.e. entanglement sudden death has been experimentally observed Brodsky_OL_2011 and theoretically studied Brodsky_PRL_2011 on the fiber-based entangled photon pair. One can enhance the PMD effect on photon pairs by increasing their differential group delays such as inserting a few polarization-maintaining fiber jumpers in their channels.
For $4\kappa\geq\gamma_{\circ}$, where the PMD effect is less dominant than the loss parameter, our system becomes non-Markovian. For simplicity, we assume $\gamma_{\circ}=4\kappa/n$ where $n>1$, then $\delta=4\kappa\sqrt{1-\frac{1}{n^{2}}}$. The probability function of the qubit is given by,
@@
Then, the harmonic function of $p_{1,2}(t)$ depends only on the PMD effect. However, our analytical approach can provide a new physical insight into the coupling constant ($\kappa$) of the PMD and the loss ($\mu$) effects on the two-photon polarization entangled state when one photon from the pair is stored in a fiber-based quantum buffer.
@@
where $E_{\circ}=(I\otimes\Gamma_{\circ})$ and $E_{1}=(I\otimes\Gamma_{1})$. The identity matrix $I=\begin{pmatrix}1&0\\ 0&1\end{pmatrix}$. The amplitude damping operators are $E_{\circ}=\begin{pmatrix}1&0\\ 0&\sqrt{1-\xi}\end{pmatrix}$ and $E_{1}=\begin{pmatrix}0&\sqrt{\xi}\\ 0&0\end{pmatrix}$, where the $\xi$ is the probability of depolarizing a photon. We measure the density matrix $\rho^{\prime}$ and obtain the density matrix elements $\rho^{\prime}_{11}=\frac{1+p^{\prime}_{11}}{4}$, $\rho^{\prime}_{22}=\frac{1-p^{\prime}_{22}}{4}$, $\rho^{\prime}_{33}=\frac{1-p^{\prime}_{33}}{4}$, $\rho^{\prime}_{44}=\frac{1+p^{\prime}_{44}}{4}$, $\rho^{\prime}_{14}=\frac{p^{\prime}_{14}}{2}$, and $\rho^{\prime}_{41}=\frac{p^{\prime}_{41}}{2}$. From each density matrix element, we write ${p^{\prime}_{nm}}_{(n,m=1,2,3,4)}$ in terms of p of the Werner State $\rho$ such as;
@@
To simplify the above expression, we assume $\Delta\phi_{h(v)}$ is relatively small and substitute $\rm{cos(\Delta\phi)}=1-\frac{\Delta\phi^{2}}{2}$ and $\rm{sin(\Delta\phi)}=\Delta\phi$ to the $\mathcal{P}_{\pm}^{f}$. The first few terms of the $\mathcal{P}_{\pm}^{f}$ are given as;
@@
This approximated probability function is depending on the power of the fiber length $L^{\frac{n}{2}}$, where the n is positive integer number.
V Data availability

### v3

@@ V Methods
Atom-Cavity-Reservoir Model. Our analytical approach is similar to a qubit of a two-level atom interacting in a high Q-cavity, which is coupled to the reservoir with the Lorentzian spectral distribution Chen_PRA_14 ; Wang_scr_17 . The cavity-reservoir coupling gives rise to the effective spectral density in the form of, $(\gamma_{\circ}\Lambda^{2})/[(\omega_{\circ}-\omega)^{2}+\Lambda^{2}]$, where the parameter $\gamma_{\circ}$ is the coupling constant between the cavity and the reservoir. The parameter $\Lambda$ is the spectral width of the reservoir. The atom transition frequency and the cavity are in resonance with $\omega_{\circ}=\omega_{\rm{cavity}}$. When $\Lambda$ becomes very large or approaches infinity, the reservoir is considered memoryless or Markovian. Then, the cavity is the main cause of the non-Markovian dynamics in the system. The diagonal (off-diagonal)density matrix elements of the atomic qubit evolve in time $t$ as a function of $p(t)\propto e^{-\frac{\gamma_{\circ}t}{2}}[cos(\frac{\delta t}{4})+\frac{\gamma_{\circ}}{\delta}sin(\frac{\delta t}{4})]^{2}$ ($\sqrt{p(t)}$), respectively, where $\delta=\sqrt{16\kappa^{2}-\gamma_{\circ}^{2}}$, and $\kappa$ is the coupling constant between the atomic qubit and the cavity and is related to the decay rate of the excited state. The parameter $\delta$ can be used to determine whether the system is Markovian ( $4\kappa<\gamma_{\circ}$) or non-Markovian ($4\kappa\geq\gamma_{\circ}$).
@@
For $4\kappa<\gamma_{\circ}$ or $1<\frac{\gamma_{\circ}}{4\kappa}$, $\delta$ becomes imaginary. The harmonic functions $sin(\cdot)$ and $cos(\cdot)$ become hyperbolic functions $cosh(\cdot)$ and $sinh(\cdot)$, respectively. Our quantum system is undergoing Markovian dynamics. This occurs when the PMD and the loss parameter are the dominant effects on the photonic qubit. Hence, the probability function $p(t)$ becomes an exponential function. The polarization mode dispersion (PMD) effect that leads to the Markovian process, i.e., entanglement sudden death, has been experimentally observed Brodsky_OL_2011 and theoretically studied Brodsky_PRL_2011 on the fiber-based entangled photon pair. One can enhance the PMD effect on photon pairs by increasing their differential group delays, such as inserting a few polarization-maintaining fiber jumpers in their channels.
For $4\kappa\geq\gamma_{\circ}$, where the PMD effect is less dominant than the loss parameter, our system becomes non-Markovian. For simplicity, we assume $\gamma_{\circ}=4\kappa/n$ where the positive real number $n>1$, then $\delta=4\kappa\sqrt{1-\frac{1}{n^{2}}}$. The probability function of the qubit is given by,
@@
Then, the harmonic function of $p_{1,2}(t)$ depends only on the PMD effect. However, our analytical approach can provide new physical insight into the coupling constant ($\kappa$) of the PMD and the loss ($\mu$) effects on the two-photon polarization entangled state when one photon from the pair is stored in a fiber-based quantum buffer.
@@
where $E_{\circ}=(I\otimes\Gamma_{\circ})$ and $E_{1}=(I\otimes\Gamma_{1})$. The identity matrix $I=\begin{pmatrix}1&0\\ 0&1\end{pmatrix}$. The amplitude damping operators are $\Gamma_{\circ}=\begin{pmatrix}1&0\\ 0&\sqrt{1-\xi}\end{pmatrix}$ and $\Gamma_{1}=\begin{pmatrix}0&\sqrt{\xi}\\ 0&0\end{pmatrix}$, where the $\xi$ is the probability of depolarizing a photon. We measure the density matrix $\rho^{\prime}$ and obtain the density matrix elements: $\rho^{\prime}_{11}=\frac{1+p^{\prime}_{11}}{4}$, $\rho^{\prime}_{22}=\frac{1-p^{\prime}_{22}}{4}$, $\rho^{\prime}_{33}=\frac{1-p^{\prime}_{33}}{4}$, $\rho^{\prime}_{44}=\frac{1+p^{\prime}_{44}}{4}$, $\rho^{\prime}_{14}=\frac{p^{\prime}_{14}}{2}$, and $\rho^{\prime}_{41}=\frac{p^{\prime}_{41}}{2}$. From each density matrix element, we write ${p^{\prime}_{nm}}_{(n,m=1,2,3,4)}$ in terms of p of the Werner State $\rho$ as:
@@
To simplify the above expression, we assume $\Delta\phi_{h(v)}$ is relatively small and substitute $\rm{cos(\Delta\phi)}=1-\frac{\Delta\phi^{2}}{2}$ and $\rm{sin(\Delta\phi)}=\Delta\phi$ into the $\mathcal{P}_{\pm}^{f}$. The first few terms of the $\mathcal{P}_{\pm}^{f}$ are given as:
@@
This approximated probability function depends on the power of the fiber length $L^{\frac{m}{2}}$, where m is a positive integer.
VI Data availability